Evaluate: (15-5)divided by [(12 divided by 2^2)-2]?
Answer by madhan_math(34)

Given (15-5)divided by [(12 divided by 2^2)-2]
consider, (12 divided by 4)-2=12/4-2=3-2=1.
Therefore, (15-5)/1=10/1=10.